OpenStudy (yacoub1993):

Solve the proportion. 10/a = 15/30 A. 2 B. 20 C. 5 D. 10

OpenStudy (anonymous):

(10)/(a)=(15)/(30) Since there is one rational expression on each side of the equation, this can be solved as a ratio. For example, (A)/(B)=(C)/(D) is equivalent to A*D=B*C. 10*30=15*a Since a is on the right-hand side of the equation, switch the sides so it is on the left-hand side of the equation. 15*a=10*30 Multiply 15 by a to get 15a. 15a=10*30 Multiply 10 by 30 to get 300. 15a=300 Divide each term in the equation by 15. (15a)/(15)=(300)/(15) Simplify the left-hand side of the equation by canceling the common factors. a=(300)/(15) Simplify the right-hand side of the equation by simplifying each term. a=20

OpenStudy (anonymous):

so, its letter B. 20

OpenStudy (yacoub1993):

thnxs bro :)

OpenStudy (anonymous):

your welcome :)

OpenStudy (anonymous):

too long.